Please help me make an equation with the given information: Focus point is (0,-1) and y=1
Answer by lwsshak3(11628)

Please help me make an equation with the given information: Focus point is (0,-1) and y=1
**
Given data shows equation is that of a parabola that opens downward.
Stand form of equation: (x-h)^2=-4p(y-k), (h,k)=(x,y) coordinates of vertex
vertex: (0,0)
p=1 (distance from focal pt to vertex)
4p=4
Equation of given parabola: x^2=-4y